# Dream to Manipulate: Compositional World Models Empowering Robot Imitation Learning with Imagination

### v1

@@ I.2 Data augmentation for Imitation learning
The recently introduced MimicGen [Mandlekar et al., 2023] demonstrated that imitation learning can benefit from data augmentation by composing different object-centric segments. The main limitation of these approaches is that they can only generate data by composing known states. In contrast, our approach generates new data by imagining novel equivariant configurations of objects and trajectories
@@
As motivated earlier, to model the object assets and their dynamics, physics engines require objects to be represented by mesh grids and not Gaussian blobs. We, therefore, adopt the 2DGS Gaussian Splatting [Huang et al., 2024] with TSDF to extract high-quality meshes, since the original Gaussian Splatting does not guarantee that Gaussians align with surfaces. Sugar [Guédon and Lepetit, 2024] proposes additional regularizations to improve mesh extraction. However, we found their meshes were not accurate enough for reliable simulations. As a physics engine, we use PyBullet as a core component.
@@
To avoid falling objects in the simulator, the table is extracted by filtering the depth by using the segmentation mask. RANSAC [Fischler and Bolles, 1981] is used to estimate a plane necessary to create a fixed convex hull functioning as a playground for the robot.
@@
For the physics simulator, the geometry of the objects is not enough; various physical parameters are also required. One can assume some physical parameters to be constant, such as the gravitational constant or the air resistance, while others are not, including the masses of objects, object elasticity, or the friction coefficient. In general, estimating physical parameters is a whole exciting field on its own [Mehra, 1974, Chebotar et al., 2019, Gao et al., 2022, Huang et al., 2023] and some of these methods, such as ASID [Memmel et al., 2024] or AdaptSim [Ren et al., 2023], can be applied to identify them. In this paper, we focus on learning a photo-realistic representation of the world powered by a physics simulator, and its usefulness for compositional manipulation world models and imitation learning. Keeping in mind the overall complexity, we consider physical parameter identification to be out of scope, yet an important direction for future work.
@@
Segmenting simulation images with Grounded-SAM [Ren et al., 2024] proved challenging due to the domain gap. While the typical gap occurs from simulation to the real world, in our case, we encountered the opposite issue. The open-vocabulary segmenter, which was primarily trained on real data, struggled to accurately detect objects, the table, and the robot in the simulation. We observed that while SAM could identify objects, grounded-DINO misclassified instances, assigning incorrect labels. Figure K.11 shows examples of predicted masks in the simple environment with a cube in the center. In contrast, using only simple prompts such as "table" and "object," and filtering out masks not on the table, the segmenter could correctly identify objects in the real data. Figure K.12 displays examples from the real setting. One possible solution to improve recognition in the simulation is to fine-tune the segmenter to align the two domains. However, since our goal is to develop a world model with real-world applicability, we chose to focus on enhancing the zero-shot capabilities for the real world.
Figure K.11: Segmentation masks in simulation. The middle column was generated with the prompt "object", whereas the right column was generated with the prompt "green cube".
Figure K.12: Examples of segmentation masks from the real images. These were generated by prompting "object" and "table".
@@
To avoid falling objects in the simulator, the table is extracted by filtering the depth by using the segmentation mask. RANSAC [Fischler and Bolles, 1981] is used to estimate a plane necessary to create a fixed convex hull. Finally, the entire scene is reconstructed, and the Gaussians of the objects are removed from the reconstruction. The procedure produced a set of Gaussians paired with the mesh for each object and the Gaussians of the scene without the extracted objects.
@@
The set of images $I$ is collected using a similar approach of Ze et al. [2023] collecting 200 images from a rotating camera. We set the resolution of the camera to $1280\times 720$, mimicking a real one. To avoid falling objects in the simulator, the table is extracted by re-projecting the depth filtered by the segmentation mask of this particular object. RANSAC [Fischler and Bolles, 1981] is used to estimate a plane, necessary to create a fixed convex hull used as a workspace. The Gaussians of each object are trained for 7000 iterations. The meshes are extracted starting from the Gaussians using TSDF and the hull. The table is used to remove the parts of the mesh penetrating the table. Without this filtering, PyBullet may detect collisions creating problems in the simulation. A similar operation should be also done in case of touching objects. The entire scene is reconstructed. The Gaussians of each object are matched with their counterpart in the entire scene using KNN and then removed to avoid duplications. The procedure produced a set of Gaussians paired with the mesh for each object, and the Gaussians of the scene without the extracted objects.
@@
Following PerAct [Shridhar et al., 2023], we use a dataset $\mathcal{D}={\zeta_{1},\zeta_{2},\ldots,\zeta_{n}}$ of $n$ demonstrations, each paired with a corresponding task encoding $\mathcal{T}={t_{1},t_{2},\ldots,t_{n}}$. For language-based goals, the task $t_{i}$ is the language embedding (e.g., a pre-trained CLIP embedding) of the phrase describing the task. Each demonstration $\zeta$ is a sequence of continuous actions $\mathcal{A}={a_{1},a_{2},\ldots,a_{t}}$ paired with observations $\mathcal{O}={\tilde{o}1,\tilde{o}2,\ldots,\tilde{o}t}$. An action $a$ consists of the 6-DoF pose, gripper open state, and whether collision avoidance was used by the motion planner to reach an intermediate pose: $a={a{\text{pose}},a{\text{open}},a{\text{collide}}}$. An observation $\tilde{o}$ consists of RGB-D images from any number of cameras. We use three cameras for simulated experiments, $\tilde{o}{\text{sim}}={o{\text{front}},o_{\text{left}},o_{\text{right}}}$, and a single camera for real-world experiments, $\tilde{o}{\text{real}}={o{\text{front}}}$. Each image in $\mathcal{O}$ has dimensions $128\times 128$.
To generate the data, the robot executes $\mathcal{A}$ in PyBullet and renders a new set of observations using simulated cameras located at the same positions as $\tilde{o}{\text{sim}}$. When the robot successfully completes $a_{i}$, the corresponding simulated observation is rendered, and the robot proceeds to execute $a{i+1}$. It is important to note that the set of cameras ${o_{\text{front}},o_{\text{left}},o_{\text{right}}}$ is not used for optimizing the Gaussian representations. Instead, the novel-view synthesis capability of Gaussian Splatting is utilized. At the end of the execution, the objects’ locations are saved. This data is then used to verify if the actions in the new environmental configurations can successfully complete the task.
@@
We noticed that the jars and shape sorter are fixed in the simulator, so we also fixed them in the real experiments. Similarly, the cupboard, which is floating in the simulator, was adjusted accordingly in the real setup. Our goal is to create a representation that the robot can manipulate and use to learn tasks that visually resemble the real-world scene.
@@
Model selection is crucial to avoid deploying suboptimal policies [Shridhar et al., 2023]. Compared to the original study, the number of validation examples was increased from 25 to 40 to enhance model selection. Similarly, the test set size was expanded from 25 to 50 to account for a broader range of evaluation scenarios. Additionally, testing was repeated five times to incorporate the variability introduced by motion planning and environmental interactions.
@@
Gaussian Splatting is designed for novel view synthesis. However, we noticed strange renderings when it is trained with high-resolution images and deployed to generate low-resolution ones, that are needed for the experiments in RLBench [James et al., 2020]. Figure L.13 shows how these problems affected the rendering of RGB and depth images. In particular, depth images must be carefully estimated since PerAct uses them to create a voxel representation of the environment. The wrong renderings in the The problem caused the robot to predict the actions incorrectly. To overcome this problem, we rendered a higher-resolution image and down-sampled it to the desired resolution.
@@
One of the most essential parts of imitation learning is data collection. At the beginning and the end of the demonstration, we recorded the images used to reconstruct the world model and checked that with the physical parameters used, the robot could replay the same trajectory. The demonstrations are collected only from the external Kinect V2 camera, positioned to resemble the front camera of RLBench as in the real experiment of PerAct [Shridhar et al., 2023]. The keypoints to execute were manually saved through the RVIZ interface in ROS1 [Quigley et al., 2009]. When the keypoints were set, we used sampling-based motion planning to generate the trajectory to execute. During the execution, we recorded a Ros bag later processed to create data in RLBench format  [^1] https://github.com/baldeeb/rosbag_data_utils/tree/main. In collecting the data, we noticed that quality is essential for correctly training the model. Sometimes formatted data contained more keypoints in similar positions, creating loops in the agent that could not exit. We encountered problems when the same task in a similar position had a different number of keypoints. We believe that these problems are reduced by increasing the batch size, but due to time constraints, we could not verify this assumption. To avoid these issues, we collected demonstrations with a constant number of keypoints, and manually cleaned the data when we noticed unwanted keypoints.

### v2

@@ I.2 Data augmentation for Imitation learning
The recently introduced MimicGen [Mandlekar et al., 2023] demonstrated that imitation learning can benefit from data augmentation by composing different object-centric segments. The main limitation of these approaches is that they can only generate data by composing known states. In contrast, our approach generates new data by imagining novel equivariant configurations of objects and trajectories.
@@
As motivated earlier, to model the object assets and their dynamics, physics engines require objects to be represented by mesh grids and not Gaussian blobs. We therefore adopt the 2DGS Gaussian Splatting [Huang et al., 2024] with TSDF to extract high-quality meshes, since the original Gaussian Splatting does not guarantee that the Gaussians align with the surfaces. Sugar [Guédon and Lepetit, 2024] proposes additional regularizations to improve mesh extraction. However, we found that their meshes were not accurate enough for reliable simulations. As a physics engine, we use PyBullet as a core component.
@@
To avoid falling objects in the simulator, the table is extracted by filtering the depth by using the segmentation mask. RANSAC [Fischler and Bolles, 1981] is used to estimate a plane necessary to create a fixed convex hull that functions as a playground for the robot.
@@
For the physics simulator, the geometry of the objects is not enough; various physical parameters are also required. Some physical parameters can be assumed to be constant, such as the gravitational constant or air resistance, while others are not, including the masses of objects, object elasticity, or the friction coefficient. In general, estimating physical parameters is a whole exciting field on its own [Mehra, 1974, Chebotar et al., 2019, Gao et al., 2022, Huang et al., 2023] and some of these methods, such as ASID [Memmel et al., 2024] or AdaptSim [Ren et al., 2023], can be applied to identify them. In this paper, we focus on learning a photo-realistic representation of the world powered by a physics simulator, and its usefulness for compositional manipulation world models and imitation learning. Keeping in mind the overall complexity, we consider physical parameter identification to be out of scope, yet an important direction for future work.
@@
Segmenting simulation images with Grounded-SAM [Ren et al., 2024] proved challenging due to the domain gap. While the typical gap occurs from simulation to the real world, in our case, we encountered the opposite issue. The open-vocabulary segmenter, which was primarily trained on real data, struggled to accurately detect objects, the table, and the robot in the simulation. We observed that while SAM could identify objects, grounded-DINO misclassified instances, assigning incorrect labels. Figure K.11 shows examples of predicted masks in the simple environment with a cube in the center. In contrast, using only simple prompts such as “table” and “object”, and filtering out masks not on the table, the segmenter could correctly identify objects in the real data. Figure K.12 displays examples from the real setting. One possible solution to improve recognition in the simulation is to fine-tune the segmenter to align the two domains. However, since our goal is to develop a world model with real-world applicability, we chose to focus on enhancing the zero-shot capabilities for the real world.
Figure K.11: Segmentation masks in simulation. The middle column was generated with the prompt “object”, whereas the right column was generated with the prompt “green cube”.
Figure K.12: Examples of segmentation masks from the real images. These were generated by prompting “object” and “table”.
@@
To avoid falling objects in the simulator, the table is extracted by filtering the depth using the segmentation mask. RANSAC [Fischler and Bolles, 1981] is used to estimate the plane necessary to create a fixed convex hull. Finally, the entire scene is reconstructed, and the Gaussians of the objects are removed from the reconstruction. The procedure produced a set of Gaussians paired with the mesh for each object and the Gaussians of the scene without the extracted objects.
@@
The set of images X is collected using a similar approach of Ze et al. [2023] collecting 200 images from a rotating camera. We set the resolution of the camera to $1280\times 720$, mimicking a real one. To avoid falling objects in the simulator, the table is extracted by re-projecting the depth filtered by the segmentation mask of this particular object. RANSAC [Fischler and Bolles, 1981] is used to estimate a plane, necessary to create a fixed convex hull used as a workspace. The Gaussians of each object are trained for 7000 iterations. The meshes are extracted starting from the Gaussians using TSDF and the hull. The table is used to remove the parts of the mesh penetrating the table. Without this filtering, PyBullet may detect collisions creating problems in the simulation. A similar operation should be also done in case of touching objects. The entire scene is reconstructed. The Gaussians of each object are matched with their counterpart in the entire scene using KNN and then removed to avoid duplications. The procedure produced a set of Gaussians paired with the mesh for each object, and the Gaussians of the scene without the extracted objects.
@@
Following PerAct [Shridhar et al., 2023], we use a dataset $\mathcal{Z}={\zeta_{1},\zeta_{2},\ldots,\zeta_{n}}$ of $n$ demonstrations, each paired with a corresponding task encoding $\mathcal{T}={t_{1},t_{2},\ldots,t_{n}}$. For language-based goals, the task $t_{i}$ is the language embedding (e.g., a pre-trained CLIP embedding) of the phrase describing the task. Each demonstration $\zeta$ is a sequence of continuous actions $\mathbf{A}={a_{1},a_{2},\ldots,a_{t}}$ paired with observations $\mathbf{Q}=(q_{1},\ldots,q_{t})$. An action $a$ consists of the 6-DoF pose, gripper open state, and whether collision avoidance was used by the motion planner to reach an intermediate pose: $a={a_{\text{pose}},a_{\text{open}},a_{\text{collide}}}$. An observation $\tilde{q}$ consists of RGB-D images from any number of cameras. We use three cameras for simulated experiments, $\tilde{q}_{\text{sim}}={q_{\text{front}},q_{\text{left}},q_{\text{right}}}$, and a single camera for real-world experiments, $\tilde{q}_{\text{real}}={q_{\text{front}}}$. Each image in $\mathcal{Q}$ has dimensions $128\times 128$.
To generate the data, the robot executes $\mathbf{A}$ in PyBullet and renders a new set of observations using simulated cameras located at the same positions as $\tilde{q}_{\text{sim}}$. When the robot successfully completes $a_{i}$, the corresponding simulated observation is rendered, and the robot proceeds to execute $a_{i+1}$. It is important to note that the set of cameras ${q_{\text{front}},q_{\text{left}},q_{\text{right}}}$ is not used for optimizing the Gaussian representations. Instead, the novel-view synthesis capability of Gaussian Splatting is utilized. At the end of the execution, the objects’ locations are saved. This data is then used to verify if the actions in the new environmental configurations can successfully complete the task.
@@
We noticed that the jars and shape sorter are fixed in the simulator, so we also fixed them in the experiments. Similarly, the cupboard, which is floating in the simulator, was adjusted accordingly in the reconstruction. Our goal is to create a representation that the robot can manipulate and use to learn tasks that visually resemble the real-world scene.
@@
Model selection is crucial to avoid deploying suboptimal policies [Shridhar et al., 2023]. Compared to the original study, the number of validation examples was increased from 25 to 40 to improve the selection of models. Similarly, the size of the test set was increased from 25 to 50 to accommodate a wider range of evaluation scenarios. In addition, the test was repeated five times to incorporate the variability introduced by motion planning and environmental interactions.
@@
Gaussian Splatting is designed for novel view synthesis. However, we noticed strange renderings when it is trained with high-resolution images and deployed to generate low-resolution ones, that are needed for the experiments in RLBench [James et al., 2020]. Figure L.13 shows how these problems affected the rendering of RGB and depth images. In particular, depth images must be carefully estimated since PerAct uses them to create a voxel representation of the environment. The rendering problem caused the robot to predict the actions incorrectly. To overcome this problem, we rendered a higher-resolution image and down-sampled it to the desired resolution.
@@
One of the most essential parts of imitation learning is data collection. At the beginning and the end of the demonstration, we recorded the images used to reconstruct the world model and checked that with the physical parameters used, the robot could replay the same trajectory. The demonstrations are collected only from the external Kinect V2 camera, positioned to resemble the front camera of RLBench as in the real experiment of PerAct [Shridhar et al., 2023]. The keypoints to execute were manually saved through the RVIZ interface in ROS1 [Quigley et al., 2009]. When the keypoints were established, we used sampling-based motion planning to generate the execution trajectory. During the execution, we recorded a Ros bag later processed to create data in RLBench format [^1] https://github.com/baldeeb/rosbag_data_utils/tree/main. In collecting the data, we noticed that quality is essential for correctly training the model. Sometimes formatted data contained more keypoints in similar positions, creating loops in the agent that could not exit. We encountered problems when the same task in a similar position had a different number of keypoints. We believe that these problems are reduced by increasing the batch size, but due to time constraints, we could not verify this assumption. To avoid these issues, we collected demonstrations with a constant number of keypoints, and manually cleaned the data when we noticed unwanted keypoints.